1)350.0 g of H2O at 30.0 oC is heated to 85.0 oC.
How many kJ of heat did it gain from the environment?
The specific heat, c, of water is 4.184 J/g·C .  

2)A sample of a metal at 90.00°C was added to 200.0 g H 2O at 21.00°C. The temperature rose to 24.50°C. [Specific heat of H 2O is 4.184 J/(g °C). Specific heat of the metal is 0.949 J/(g°C). How many grams of metal was it?\

3)Use the information in this thermochemical equation to solve the following problem.   
AgNO3   +   HCl   -> AgCl + HNO3     
75.0 mL of 0.840 M AgNO3 and 75.0 mL of 0.840 M HCl are mixed in a constant-pressure (coffee-cup) calorimeter.   The two solutions are both initially at 20.00 degree centigrade. Final temperature of combined solution is 26.80 oC. Assume that the aqueous solutions have the same specific heat as water (4.184 J/ g oC ) and the same density of water (1.00 g/mL) .
  
How much heat did the reaction release?
  
Hint: no limiting reagent, substances are presented in correct molar ratios.
Answer #1

1)
Given:
m = 350 g
C = 4.184 J/g.oC
Ti = 30 oC
Tf = 85 oC
use:
Q = m*C*(Tf-Ti)
Q = 350.0*4.184*(85.0-30.0)
Q = 8.05*10^4 J
Q = 80.5 KJ
Answer: 80.5 KJ

Only 1 question at a time please